OpenStudy (woahning):

I'll give you a medal and become your fan. This is the last question and I'm just terrible at this topic. Is the line through points P(-8,-10) and Q(-5.-12) perpendicular to the line through points R(9,-6) and S(17,-5)? Explain?

OpenStudy (perl):

compare the slopes of the two lines

OpenStudy (perl):

slope PQ = -2/3 slope RS = 1/8

OpenStudy (perl):

the lines are neither parallel (same slope) nor perpindicular ( negative reciprocal )

OpenStudy (woahning):

is it like the y2-y1/x2-x1 formula

OpenStudy (perl):

yes

OpenStudy (perl):

that is how i calculated the slope

OpenStudy (woahning):

okay thank you so much